# Axion Astrophysics

### v1

@@ 7 Supernovae and Neutron Stars
Core-collapse SNe correspond to the terminal phase of massive stars (with mass $M\gtrsim 8$ M), which become unstable at the end of their lives, collapsing and ejecting their outer mantle in a shock-wave drive explosion. During such a process $99~{}\%$ of the emitted energy [$E\sim{\mathcal{O}}(10^{53})$ erg] is released by neutrinos and antineutrinos of all the flavors with average energies of 15 MeV, on a timescale of ten seconds. Due to the huge amount of energy released in neutrinos, a SN can be really considered a cosmic neutrino factory. During such an event, neutrinos play the role of astrophysical messengers allowing to probe both fundamental neutrino properties (like mass, mixing, interactions) and the SN explosion mechanism (see, e.g. Refs. [448, 449, 450, 451, 452, 453] for reviews). Furthermore, SNe are powerful laboratories to study novel weakly interacting light particles (see, e.g., Refs. [45, 454, 15]). For typical SN core temperatures, $T\sim{\mathcal{O}}(30)$ MeV, particles with masses $m_{a}<100$ MeV can be thermally produced without being Boltzmann suppressed.
@@
In the specific case of axions, different arguments have been proposed to constrain their emissivity (see Refs. [15, 46] for reviews). If sufficiently weakly coupled, axions stream freely (i.e., without re-interacting with the stellar medium) out of the star, providing a very efficient cooling mechanism. This consideration has allowed to set stringent bounds on the axion-nucleon coupling from the comparison with the observation of the SN 1987A neutrino burst (see, e.g., Refs. [464, 168, 465, 169]). This argument, however, cannot be used to constrain arbitrarily large couplings. In fact, for a sufficiently large coupling, axions would be trapped in the SN core and thermalize. In this case, axions would be emitted from a surface, the axion-sphere, in close analogy with the neutrino case. Even trapped axions may therefore proved an efficient energy-transfer mechanism which leads to further constraints [464, 170]. This topic was revisited from a modern perspective in recent years, as we will document in the Section. Furthermore, in the case of massive axions (with mass $m_{a}\gtrsim 10$ MeV) coupled with photons one can have also the possibility that these particles decay into photons into the star behind the shock-wave, leading to an energy-deposition which can be used to obtain other constraints requiring that this effect does not lead to too energetic explosions [474]. Conversely, if the axion-photon coupling is sufficiently small, the decay can happen outside the SN leading to an observable gamma-ray signal. A similar signal is expected also in the case of ultra-light axions (with $m_{a}\lesssim 10^{-10}$ eV) converting into photons in the Galactic magnetic-field [475, 476, 297, 339]. Therefore, one expects a lot of opportunities from a SN explosion to probe axions as we will document in details in this Section.
In Sec. 7.1 we will recall the main steps leading to a core-collapse SN explosion. In Sec. 7.2 we present the expected features of the SN neutrino burst and we compare them with SN 1987A neutrino observation. In Sec. 7.3 we discuss how the duration of the SN 1987A neutrino burst can be used as a tool to constrain FIP properties. In particular, we discuss the cooling bound both the case of free-streaming and trapped particles. In Sec. 7.4 we focus on axions coupled with nucleons, discussing in details the emissivity in nuclear medium and the progresses in the SN mass bound. Then, in Sec. 7.5 we consider axions coupled with photons and we discuss the bounds coming gamma-ray signal and energy-deposition. in Sec. 7.6 we comment on other axion scenarios recently discussed in literature. Finally, in Sec. 7.7 the impact of QCD axion on neutron star cooling is also reviewed.
@@ 7.1 SN explosion mechanism
    Initial phase of the collapse (post-bounce time $t_{\rm pb}\sim 0$): The collapse of the core starts when the temperature is high enough to dissociate iron nuclei by means of photodisintegration $\gamma~\,{}^{56}{\rm Fe}\rightarrow 13\alpha~{}4n$. At the beginning of the collapse, the electron degeneracy is reduced due to the electron capture $ep\rightarrow n\nu_{e}$ and neutrinos freely escape. The collapse is accelerated by the lower degeneracy pressure, heavier nuclei fall in the inner star, where their $\beta$ decay subtracts energy from the core. This is the first neutronization phase.
@@
The scenario depicted above is called the neutrino-driven mechanism, since essentially an explosion is associated with the shock-wave reheating by neutrinos. The main challenge of SN simulations is to realize self-consistently this complicated dynamics [486, 487]. Indeed, one of the main problems faced by SN simulations is related to the fact that often the shock-wave cannot be self-consistently revitalized by the neutrino-energy deposition. This is a typical problem often encountered in one-dimensional spherically symmetric SN models (except for low-mass electron capture SNe, with masses in the range $8-10$ $M_{\odot}$ [488, 489]). In this situation an explosion can be triggered artificially enhancing the neutrino energy deposition behind the shock-wave. However, it is expected that in multi-dimensional SN simulations [490, 491, 492, 493, 494] the explosion would be naturally achieved. Indeed, the onset of the explosion would be supported by matter convective motions helped by hydrodynamical instabilities [495, 496] such as the Standing Accretion Shock Instability (SASI) [497] which enhances the neutrino-energy deposition. At this regard, recent two-dimensional SN models are successfully exploding (see, e.g., Ref. [498] for an inspection of different models). The forefront of the research has reached the development of exploding three-dimensional SN simulations by different groups (see, e.g., Refs. [499, 500, 501, 490, 491]). Despite these significant progresses, a proof of the robustness of this neutrino heating scenario is still lacking. Alternative scenarios have been also proposed involving magnetohydrodynamics [502], acoustics [503] or phase-transition mechanisms [504]. However, for definitiveness in the following we will refer only to the neutrino-heating picture described above. Indeed, only in this framework the SN axion emission has been characterized. Furthermore, since in order to constrain axions we will be interested in the SN cooling phase up to $t\sim 10-20$ s after core-bounce, we will consider only 1D models which are currently the only ones extending till such late times (see, e.g., Ref. [463]).
@@ 7.4 Axion-nucleon coupling
In general, in the case of canonical QCD axion models, the SN arguments rule out QCD axion masses $m_{a}\gtrsim\mathcal{O}(10)~{}{\rm meV}$. This is in contrast with the original literature on the SN 1987A bound, which reported the existence of a window around a QCD axion mass $m_{a}\sim{\mathcal{O}}(1)$ eV, classicaly dubbed the “hadronic axion window” [535]. Finally, this bound is stronger than the reach of current and future cosmological experiments, which would probe axion masses $m_{a}\gtrsim 150~{}{\rm meV}$ [536]. Thus, it is unlikely that future cosmological probes would find signatures of the QCD axion mass as hot dark matter. Nevertheless, lighter axion can still play the role of dark radiation (see, e.g., Ref. [537]). Recently, in Ref. [538] it has been performed a quantitative study of the role of SN axion emission from hadronic matter beyond the first generation–in particular strange matter, obtaining a bound on the axial axion-strange-strange coupling, as well as on the axion-down-strange counterpart. In particular, this latter coupling can be as small as $10^{-2}$ for $f_{a}\sim 10^{9}$ GeV.
@@ 7.5.2 Massive axions
Figure 57: Constraints on the axion-photon coupling $g_{a\gamma}$ as a function of mass $m_{a}$ in the high mass regime. The gray region labeled ‘SN 1987A ($\nu$)’ is the cooling bound on axions coupled to photons [41]. The red exclusion region labeled ‘SN 1987A’ derives from gamma-ray decays of axions produced in SN 1987A and measured by SMM [338]. The red regions labeled ‘SN 2023ixf’ [547] derives from similar arguments applied to SN 2023ixf and Fermi-LAT observations (magenta band). The other, semi-transparent regions are constraints from the energy axions could deposit in low-energy SNe (orange) and the diffuse gamma-ray flux due to decays of axions produced in all past SNe (yellow), both from Ref. [474]; changes in the evolution of horizontal branch stars (purple) [33, 147, 151]; the non-observation of X-rays after the multi-messenger observation GW170817 of a neutron star merger [548] (gray); the irreducible cosmic axion density from freeze-in production (blue) [295]; as well as from the dissociation of light elements during BBN (green) [549], the most conservative bound. Figure adapted from Ref. [547]. (Courtesy of Eike Ravensburg).
@@ 7.6 Beyond the minimal scenario
Apart from the case of axions coupled with nucleons and photons, the SN bound has been studied for different types of axions with a variety of couplings. We give here a brief overview of different cases explored in the literature. Refs. [565, 566] studied the case of axions coupled with electrons, which would be produced in a SN via electron-proton bremsstrahlung and electron-positron fusion, possibly including loop effects. In this case, the SN 1987A cooling bound would exclude currently unprobed regions down to $g_{ae}\sim 2.5\times 10^{-10}$ at $m_{a}\sim 120$ MeV. In Ref. [528] (see also Ref. [567]) it was considered the case of muonphilic bosons both scalars $\phi$ or pseudoscalars $a$ that couple to muons through the Yukawa operators
@@
In Ref. [196] (see also Ref. [190]) were considered axions coupled to nucleons and photons only through the nucleon electric-dipole moment portal (see Sec. 3.2.4). As discussed before, this coupling is a model-independent feature of QCD axions, which solve the strong CP problem, and might arise as well in more general axion-like particle setups. The SN 1987A bound allows the exclusion of the coupling to EDM in the range $6.7\times 10^{-9}{\rm GeV}^{-2}\lesssim g_{d}\lesssim 7.7\times 10^{-6}{\rm GeV}^{-2}$. Intriguingly, there is potential to detect the axion flux through a distinctive gamma-ray signal generated in a large underground neutrino detector, such as the proposed Hyper-Kamiokande, even for couplings below the current lower bound.
@@ 7.7 Neutron stars cooling
Older NSs, with ages $\sim 10^{5}$$10^{6}$ yrs, have also been studied. In particular, Refs. [585, 586] considered the expected X-ray signal deriving from the conversion of the NS axion flux in the magnetic fields surrounding the NS. Furthermore, the analysis in Ref. [586] found that the observed excess of hard $X$-ray emission in the $2-8$ keV energy range from the nearby Magnificent Seven isolated NSs could be explained by this conversion of NS axions, if axions have mass $m_{a}\lesssim 2\times 10^{-5}$ eV and $g_{a\gamma}\times g_{an}\in(2\times 10^{-21},10^{-18})$ GeV-1 at 95% C.L.. Some attempts to present this as QCD axions were also considered in the literature [139]. It is finally worth mentioning the case of NS merger [587], which however finds less restrictive bounds with respect to the ones obtained from SNe and NS.
@@ 8 Discussions and Conclusions
What will we learn in case of a direct detection of an axion signal from the stars? Apart from cornering the axion mass and coupling with photons, axions will constitute a new probe to learn about the inner structure of stars, complementary with the information that one can get from photons and neutrinos.
@@
What are the prospects for improving the energy-loss argument? We feel quite optimistic in this respect. An unparalleled progress in the asteroseismology of white dwarfs is expected in the coming years, thanks to space missions such as TESS, and in globular clusters due to the data releases from the astrometric satellite Gaia. Regarding the supernova and neutron star environments, improving the cooling bounds will require better characterization of axion emissivity processes in dense nuclear matter. In recent years, several studies have been conducted addressing, for example, corrections beyond the one-pion approximation and the inclusion of pionic processes. However, further research is necessary to achieve a coherent understanding. Moreover, the SN 1987A bound is based on the observation of a sparse neutrino signal. New state-of-the-art supernova neutrino simulations, including axions, are essential to assess the true impact of axion energy loss on the neutrino signal and to interpret the high-statistics neutrino data from the next Galactic supernova.
@@
In conclusion, with this review, we have presented the state of the art in stellar bounds and opportunities to study axions. We hope this work will be useful to the community in guiding future theoretical and experimental studies on axions and will stimulate new ideas and further investigations to improve the arguments discussed here.
@@ Acknowledgments
We warmly thank Alessandro Lella and Giuseppe Lucente for the many disussions during the development of this work. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@
MG acknowledges support from the Spanish Agencia Estatal de Investigación under grant PID2019-108122GB-C31, funded by MCIN/AEI/10.13039/501100011033, and from the “European Union NextGenerationEU/PRTR” (Planes complementarios, Programa de Astrofísica y Física de Altas Energías). He also acknowledges support from grant PGC2022-126078NB-C21, “Aún más allá de los modelos estándar,” funded by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe.” Additionally, MG acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the European Research Council (ERC) grant agreement ERC-2017-AdG788781 (IAXO+) and OS acknowledges funding from the large program BRAVOSUN of the Italian National Institute of Astrophysics.

### v2

@@ 7 Supernovae and Neutron Stars
Core-collapse SNe correspond to the terminal phase of massive stars (with mass $M\gtrsim 8$ M), which become unstable at the end of their lives, collapsing and ejecting their outer mantle in a shock-wave driven explosion. During such a process $99~{}\%$ of the emitted energy [$E\sim{\mathcal{O}}(10^{53})$ erg] is released by neutrinos and antineutrinos of all the flavors with average energies of 15 MeV, on a timescale of ten seconds. Due to the huge amount of energy released in neutrinos, a SN can be really considered a cosmic neutrino factory. During such an event, neutrinos play the role of astrophysical messengers allowing us to probe both fundamental neutrino properties (like mass, mixing, interactions) and the SN explosion mechanism (see, e.g. Refs. [461, 462, 463, 464, 465, 466] for reviews). Furthermore, SNe are powerful laboratories to study novel weakly interacting light particles (see, e.g., Refs. [45, 467, 15]). For typical SN core temperatures, $T\sim{\mathcal{O}}(30)$ MeV, particles with masses $m_{a}<100$ MeV can be thermally produced without being Boltzmann suppressed.
@@
In the specific case of axions, different arguments have been proposed to constrain their emissivity (see Refs. [15, 46] for reviews). If sufficiently weakly coupled, axions stream freely (i.e., without re-interacting with the stellar medium) out of the star, providing a very efficient cooling mechanism. This consideration has allowed to set stringent bounds on the axion-nucleon coupling from the comparison with the observation of the SN 1987A neutrino burst (see, e.g., Refs. [477, 170, 478, 171]). This argument, however, cannot be used to constrain arbitrarily large couplings. In fact, for a sufficiently large coupling, axions would be trapped in the SN core and thermalize. In this case, axions would be emitted from a surface, the axion-sphere, in close analogy with the neutrino case. Even trapped axions may therefore provide an efficient energy-transfer mechanism which leads to further constraints [477, 172]. This topic was revisited from a modern perspective in recent years, as we will document in this Section. Furthermore, in the case of massive axions (with mass $m_{a}\gtrsim 10$ MeV) coupled with photons one can have also the possibility that these particles decay into photons into the star behind the shock-wave, leading to an energy-deposition which can be used to obtain other constraints requiring that this effect does not lead to too energetic explosions [487]. Conversely, if the axion-photon coupling is sufficiently small, the decay can happen outside the SN leading to an observable gamma-ray signal. A similar signal is expected also in the case of ultra-light axions (with $m_{a}\lesssim 10^{-10}$ eV) converting into photons in the Galactic magnetic-field [488, 489, 300, 346]. Therefore, one expects a lot of opportunities from a SN explosion to probe axions as we will document in detail in this Section.
In Sec. 7.1 we will recall the main steps leading to a core-collapse SN explosion. In Sec. 7.2 we present the expected features of the SN neutrino burst and we compare them with the SN 1987A neutrino observation. In Sec. 7.3 we discuss how the duration of the SN 1987A neutrino burst can be used as a tool to constrain FIP properties. In particular, we discuss the cooling bound in both the case of free-streaming and trapped particles. In Sec. 7.4 we focus on axions coupled with nucleons, discussing in detail the emissivity in nuclear medium and the progresses in the SN mass bound. Then, in Sec. 7.5 we consider axions coupled with photons and we discuss the bounds coming from gamma-ray signal and energy-deposition. in Sec. 7.6 we comment on other axion scenarios recently discussed in the literature. Finally, in Sec. 7.7 the impact of QCD axions on neutron star cooling is also reviewed.
@@ 7.1 SN explosion mechanism
    Initial phase of the collapse (post-bounce time $t_{\rm pb}\sim 0$): The collapse of the core starts when the temperature is high enough to dissociate iron nuclei by means of photodisintegration $\gamma~\,{}^{56}{\rm Fe}\rightarrow 13\alpha~{}4n$. At the beginning of the collapse, the electron degeneracy is reduced due to the electron capture $ep\rightarrow n\nu_{e}$ and neutrinos freely escape. The collapse is accelerated by the lower degeneracy pressure, causing heavier nuclei to fall in the inner star, where their $\beta$ decay subtracts energy from the core. This is the first neutronization phase.
@@
The scenario depicted above is called the neutrino-driven mechanism, since essentially an explosion is associated with the shock-wave reheating by neutrinos. The main challenge of SN simulations is to realize self-consistently this complicated dynamics [499, 500]. Indeed, one of the main problems faced by SN simulations is related to the fact that often the shock-wave cannot be self-consistently revitalized by the neutrino-energy deposition. This is a typical problem often encountered in one-dimensional spherically symmetric SN models (except for low-mass electron capture SNe, with masses in the range $8-10$ $M_{\odot}$ [501, 502]). In this situation an explosion can be triggered artificially enhancing the neutrino energy deposition behind the shock-wave. However, it is expected that in multi-dimensional SN simulations [503, 504, 505, 506, 507] the explosion would be naturally achieved. Indeed, the onset of the explosion would be supported by matter convective motions helped by hydrodynamical instabilities [508, 509] such as the Standing Accretion Shock Instability (SASI) [510] which enhances the neutrino-energy deposition. At this regard, recent two-dimensional SN models are successfully exploding (see, e.g., Ref. [511] for an inspection of different models). The forefront of the research has reached the development of exploding three-dimensional SN simulations by different groups (see, e.g., Refs. [512, 513, 514, 503, 504]). Despite these significant advances, a proof of the robustness of this neutrino heating scenario is still lacking. Alternative scenarios have been also proposed involving magnetohydrodynamics [515], acoustics [516] or phase-transition mechanisms [517]. However, for definitiveness in the following we will refer only to the neutrino-heating picture described above. Indeed, only in this framework the SN axion emission has been characterized. Furthermore, since in order to constrain axions we will be interested in the SN cooling phase up to $t\sim 10-20$ s after core-bounce, we will consider only 1D models which are currently the only ones extending till such late times (see, e.g., Ref. [476]).
@@ 7.4 Axion-nucleon coupling
In general, in the case of canonical QCD axion models, the SN arguments rule out QCD axion masses $m_{a}\gtrsim\mathcal{O}(10)~{}{\rm meV}$. This is in contrast with the original literature on the SN 1987A bound, which reported the existence of a window around a QCD axion mass $m_{a}\sim{\mathcal{O}}(1)$ eV, classicaly dubbed the “hadronic axion window” [548]. Finally, this bound is stronger than the reach of current and future cosmological experiments, which would probe axion masses $m_{a}\gtrsim 150~{}{\rm meV}$ [549]. Thus, it is unlikely that future cosmological probes would find signatures of the QCD axion mass as hot dark matter. Nevertheless, lighter axion can still play the role of dark radiation (see, e.g., Ref. [550, 551, 552]). Recently, in Ref. [553] it has been performed a quantitative study of the role of SN axion emission from hadronic matter beyond the first generation–in particular strange matter, obtaining a bound on the axial axion-strange-strange coupling, as well as on the axion-down-strange counterpart. In particular, this latter coupling can be as small as $10^{-2}$ for $f_{a}\sim 10^{9}$ GeV.
@@ 7.5.2 Massive axions
Figure 57: Constraints on the axion-photon coupling $g_{a\gamma}$ as a function of mass $m_{a}$ in the high mass regime. The gray region labeled ‘SN 1987A ($\nu$)’ is the cooling bound on axions coupled to photons [41]. The red exclusion region labeled ‘SN 1987A’ derives from gamma-ray decays of axions produced in SN 1987A and measured by SMM [345]. The red regions labeled ‘SN 2023ixf’ [562] derives from similar arguments applied to SN 2023ixf and Fermi-LAT observations (magenta band). The other, semi-transparent regions are constraints from the energy axions could deposit in low-energy SNe (orange) and the diffuse gamma-ray flux due to decays of axions produced in all past SNe (yellow), both from Ref. [487]; changes in the evolution of horizontal branch stars (purple) [33, 148, 153]; the non-observation of X-rays after the multi-messenger observation GW170817 of a neutron star merger [563, 564] (gray); the irreducible cosmic axion density from freeze-in production (blue) [298]; as well as from the dissociation of light elements during BBN (green) [565], the most conservative bound. Figure adapted from Ref. [562]. (Courtesy of Eike Ravensburg).
@@ 7.6 Beyond the minimal scenario
Apart from the case of axions coupled with nucleons and photons, the SN bound has been studied for different types of axions with a variety of couplings. We give here a brief overview of different cases explored in the literature. Refs. [582, 583] studied the case of axions coupled with electrons, which would be produced in a SN via electron-proton bremsstrahlung and electron-positron fusion, possibly including loop effects. In this case, the SN 1987A cooling bound would exclude currently unprobed regions down to $g_{ae}\sim 2.5\times 10^{-10}$ at $m_{a}\sim 120$ MeV. In Ref. [541] (see also Ref. [584]) it was considered the case of muonphilic bosons both scalars $\phi$ or pseudoscalars $a$ that couple to muons through the Yukawa operators (see Ref. [585] for light particles coupled to tau leptons)
@@
In Ref. [198] (see also Ref. [192]) were considered axions coupled to nucleons and photons only through the nucleon electric-dipole moment portal (see Sec. 3.2.4). As discussed before, this coupling is a model-independent feature of QCD axions, which solve the strong CP problem, and might arise as well in more general axion-like particle setups. [^32] More recently, two unrelated model-independent SN axion bounds have been proposed in Ref. [591] and in Ref. [555]. These bounds are subdominant for most QCD axion models, which typically also couple to fermions. However, while the coupling to fermions can, in principle, be suppressed, the bounds derived in the aforementioned references—similar to those derived from the nEDM—apply universally to all QCD axion models. The SN 1987A bound allows the exclusion of the coupling to EDM in the range $6.7\times 10^{-9}{\rm GeV}^{-2}\lesssim g_{d}\lesssim 7.7\times 10^{-6}{\rm GeV}^{-2}$. Intriguingly, there is potential to detect the axion flux through a distinctive gamma-ray signal generated in a large underground neutrino detector, such as the proposed Hyper-Kamiokande, even for couplings below the current lower bound.
@@ 7.7 Neutron stars cooling
Older NSs, with ages $\sim 10^{5}$$10^{6}$ yrs, have also been studied. In particular, Refs. [604, 605] considered the expected X-ray signal deriving from the conversion of the NS axion flux in the magnetic fields surrounding the NS (see also Refs. [606, 607]). Furthermore, the analysis in Ref. [605] found that the observed excess of hard $X$-ray emission in the $2-8$ keV energy range from the nearby Magnificent Seven isolated NSs could be explained by this conversion of NS axions, if axions have mass $m_{a}\lesssim 2\times 10^{-5}$ eV and $g_{a\gamma}\times g_{an}\in(2\times 10^{-21},10^{-18})$ GeV-1 at 95% C.L.. Some attempts to present this as QCD axions were also considered in the literature [140]. It is finally worth mentioning the case of NS merger [608], which however finds less restrictive bounds with respect to the ones obtained from SNe and NS.
@@ 8 Discussions and Conclusions
What will we learn in the case of a direct detection of an axion signal from the stars? Apart from cornering the axion mass and coupling with photons, axions will constitute a new probe to learn about the inner structure of stars, complementary with the information that one can get from photons and neutrinos.
@@
What are the prospects for improving the energy-loss argument? We feel quite optimistic in this respect. An unparalleled progress in the asteroseismology of white dwarfs is expected in the coming years, thanks to space missions such as TESS, and in globular clusters due to the data releases from the astrometric satellite Gaia. Regarding the supernova and neutron star environments, improving the cooling bounds will require better characterization of axion emissivity processes in dense nuclear matter. In recent years, several studies have been conducted addressing, for example, corrections beyond the one-pion approximation and the inclusion of pionic processes. However, further research is necessary to achieve a coherent understanding. Moreover, the SN 1987A bound is based on the observation of a sparse neutrino signal. New state-of-the-art 3D supernova neutrino simulations [503, 504], including axions, are essential to assess the true impact of axion energy loss on the neutrino signal and to interpret the high-statistics neutrino data from the next Galactic supernova.
@@
In conclusion, with this review, we have presented the state-of-the-art in stellar bounds and opportunities to study axions. We hope this work will be useful to the community in guiding future theoretical and experimental studies on axions and will stimulate new ideas and further investigations to improve the arguments discussed here.
@@ Acknowledgments
We warmly thank Alessandro Lella and Giuseppe Lucente for the many discussions during the development of this work, and Georg Raffelt for several relevant comments and suggestions. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@
MG acknowledges support from the Spanish Agencia Estatal de Investigación under grant PID2019-108122GB-C31, funded by MCIN/AEI/10.13039/501100011033, and from the “European Union NextGenerationEU/PRTR” (Planes complementarios, Programa de Astrofísica y Física de Altas Energías). He also acknowledges support from grant PGC2022-126078NB-C21, “Aún más allá de los modelos estándar,” funded by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe.” Additionally, MG acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the European Research Council (ERC) grant agreement ERC-2017-AdG788781 (IAXO+).
OS acknowledges funding from the large program BRAVOSUN of the Italian National Institute of Astrophysics.